4. The sum of all solutions of the equation 2(2x – 1) -+ (2x - 1) = 1 is A) B) D) 9 NIN c) E) 2 21 16
Answer #1

-213 2 (2X -1) + (2x-1)=173 Let (2x-1)3 = D then 2 A2 + من =0 A = 1 2 AP + A-1 2 A² + 2A - A-1 2 A CA+) -1 (Atl) = 0 (2A-) (A2x -1 -1 = 2x = - 1 tl A r X = o & if * A= (2x -1)-Y ñ - unt o 2X-1 = (1) 2 x - 1 8 2 X 8+ X = 9 코sum of all roots 9 +0 o 이 이 2